Can someone help me understand this

You develop the following information. Your firm has a target capital structure of 70% common equity, 5% preferred stock and 25% debt. The firms tax rate is 25%. The firm can issue up to $200,000 worth of debt at a before-tax cost of 9%. Then it will cost the firm 11% before-tax on debt up to $400,000 After that point, the before-tax cost of debt will be 13%. The firms preferred stock carries an annual dividend of $2 per share The issue price of the preferred would be $25 with 1.5% of the issue price charged as flotation costs. The firm expects to have $950,000 in earnings and have a dividend payout ratio of 65%. The firm bases its cost of retained earnings on the CAPM approach. For this purpose, you determine the growth rate of the market will be 5% and the market dividend yield is 2%. The risk-free rate is 3%. The firms beta is 1.05 The firm can issue new common stock with a S0.50 dividend, price of $20 per share, flotation costs of 1.75% of issue price and growth rate expected of 6%. This holds for up to $630,000 in equity after which it will cost 10% for new common stock. 1) Determine the marginal cost of capital schedule. 2) Determine which projects you would accept and what the optimal capital budget would be by combining the investment opportunity schedule (information below) and marginal cost of capital schedule Iniial Cost $375,000 $300,000 $175,000 $100,000 $200,000 IRR 8.5% 11% 10% 7.5% 6% Projec
Answer #1

Soln : Here we need to calculate the Marginal cost of capital = weighted average *costof (debt + preferred stock + common equiy)

Step 1 : Rate of return of equity to be calculated :

Using CAPM approach , return, r = risk free rate + market premium * beta = 3 + 1.05*(5+2-3) = 3+ 1.05*4 = 7.20%

Also, dividend on equity are also to be paid i.e. = 0.50/20 = 2.5% and cost of equity on new issue would be = 2.5 +1.75 + 6 = 10.25% = cost of equity taken for new issue

Cost of preffered stock = 2/25 = 8%

weight of preffered stock = 5%, weight of common equity = 70%

cost of debt = 9% upto 200000, after that 11% upto 400000

After tax cost of debt = 9%*(1-25%) = 9*75% = 6.75%

Marginal cost of capital schedule = 25%* 6.75 + 5%*8 +70%* 10.25 = 9.26%

2) Project C will be the better option as it is closest to the marginal cost of capital and intersecting of investment opportunity to marginal cost of capital is happening here that will be optimal choice.